Circular No. 7366

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000H IN IC 454 P. Garnavich, University of Notre Dame; and P. Challis, S. Jha, and

R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 2000H was obtained by K. Rines on Feb. 9.2 UT with the F. L. Whipple Observatory 1.5-m telescope, showing it to be a type-Ia supernova. The characteristic Si II 615-nm absorption is present, but with a pronounced asymmetric profile, being very steep on the red side of the line. Assuming a recession velocity of 4170 km/s, based on the narrow emission from the host galaxy, the photospheric velocity from the Si II absorption minimum is only −4100 km/s. The 580-nm feature is deep relative to the Si II, which suggests that this is a subluminous, fast-declining event. The peak observed magnitude of 17 (IAUC 7361) also suggests a very low luminosity for a type-Ia event.

XTE J1543–568 M. H. Finger, Universities Space Research Association, and C. A. Wil-

son, Marshall Space Flight Center, NASA, report for the Compton Gamma Ray Observatory (CGRO) BATSE team: “Pulsations from the newly discovered x-ray pulsar XTE J1543–568 (IAUC 7363) are being detected in the hard x-ray band with BATSE. Detectable pulses are found in BATSE data beginning on Jan. 25; from the data of Jan. 25.0–Feb. 2.0 UT, a frequency of 36.86162(6) mHz at Jan. 29.0 and a frequency rate of 1.41(6) × 10^{−11} Hz/s were determined. The pulse profile is double-peaked, with one peak being less than half the height of the other. The pulsed flux (mean minimum) for this interval was 11 ± 3 mCrab (20–50 keV), and the mean flux measured using earth occultations was 32 ± 12 mCrab (20–50 keV). The pulse frequency increased to 36.87211(14) mHz by Feb. 8.0, and the pulsed flux to 20 ± 4 mCrab, as determined from the data of Feb. 6.0–8.0. No occultation fluxes were available in this interval because of the CGRO spacecraft orbit.”

NOVA SAGITTARII 2000 Visual magnitude estimates by A. Pearce, Nedlands, W. Australia:

Feb. 12.829 UT, 12.0; 13.819, 12.4.

2000 February 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green